OpenStudy (anonymous):

using the squar box formula waht is (x+4)^2

OpenStudy (amistre64):

square box formula? can you state the formula?

OpenStudy (anonymous):

Its the square box equation?

OpenStudy (amistre64):

i think this is what they mean

OpenStudy (amistre64):

which amounts to this: x^2 +8x +16

OpenStudy (anonymous):

\[\left[ + \right] .\] it looks somthing like this, it is used to find the product

OpenStudy (anonymous):

square box formula! ask your teacher if a square lives in two dimensions or three.

OpenStudy (anonymous):

we are just being silly. \[(a+b)^2=a^2+2ab+b^2\] so you get just what amistre said.

OpenStudy (anonymous):

im not in mty math class but the equation is (x+4)^2

OpenStudy (anonymous):

I looked on google too.

OpenStudy (anonymous):

\[(x+4)^2=(x+4)(x+4)=x^2+4x+4x+16=x^2+8x+16\]

OpenStudy (anonymous):

Thank you(: